OpenStudy (anonymous):

Find the measures of the angles. A. m∠x B. m∠y ---------- A. m∠x = 50° m∠y = 135° B. m∠x = 125° m∠y = 50° C. m∠x = 50° m∠y = 125° D. m∠x = 135° m∠y = 50°

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

did you still need help? or were you able to figure it out?

OpenStudy (anonymous):

I still need help!

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you agree that angle ADE is a right angle?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so ADE is 90 degrees we see that a piece of ADE is 40 degrees. The other piece is x so... x+40 = 90 what must x be?

OpenStudy (anonymous):

50

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

as for the angle with y in it, the 55 degree angle and the angle y add up to 180 degrees because these two angles form a straight angle (which is 180 degrees) so 55+y = 180

OpenStudy (anonymous):

125

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

Can you help with another?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

What are the values of x and y in the diagram? A. x = 77° y = 93 B. x = 87° y = 83° C. x = 77° y = 83° D. x = 87° y = 93°

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

focus on the triangle Rule: in any triangle, the 3 angles always add to 180 degrees

jimthompson5910 (jim_thompson5910):

so x+67+36 = 180